A $26,000 loan borrowed 5 years ago, has been scheduled for re-pay starting today. The payment schedule is equal monthly payments made at the end of every month for the next 10 years. Find the size of the payments if interest is 3.6% p.a. compounding monthly
Answer #1
Value of loan today
FV= PV*(1+r)^n
Where,
FV= Future Value
PV = Present Value
r = Interest rate =3.6%/12 =0.30%
n= periods in number =12*5 =60
= $26000*( 1+0.003)^60
=26000*1.19689
= $31119.26
EMI = [P x R x (1+R)^N]/[(1+R)^N-1]
Where,
EMI= Equal Monthly Payment
P= Loan Amount
R= Interest rate per period =0.3%
N= Number of periods =12*10 =120
= [ $31119.26x0.003 x (1+0.003)^120]/[(1+0.003)^120 -1]
= [ $93.35778( 1.003 )^120] / [(1.003 )^120 -1
=$309.19
Size of payment =$309.19